OpenStudy (anonymous):

On Monday, a photographer took 16 portraits and 12 of them were of children. On Tuesday, he took 24 portraits and 20 of them were of children. Determine if the same fraction of portraits taken each day were of children.

OpenStudy (anonymous):

A. Yes, the fractions of portraits that were children are the same. B. No, the fractions of portraits that were children are not the same.

OpenStudy (anonymous):

@LegendarySadist @iGreenHacked

OpenStudy (anonymous):

Just simplify \[\large \sf \frac{12}{16}\] and \[\large \sf \frac{20}{24}\]

OpenStudy (anonymous):

Okay so it is 3/4 and 5/6

OpenStudy (anonymous):

Does 3/4 = 5/6?

OpenStudy (anonymous):

No so the answer is B?

OpenStudy (anonymous):

Correct, it would be B

OpenStudy (anonymous):

Thanks!!!!